# Singleton Park, Swansea SA2 8PP, UK

### v1

@@ Abstract
Following Polchinski and Strassler Polchinski:2001tt and our previous work Armoni:2024gqv, we study high-energy pion scattering in the holographic QCD hard-wall model.  In particular, we focus on comparing our predictions for the angular dependence of $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering with experimental data extracted from the process $\pi^{-}p\to\pi^{+}\pi^{-}n$.  Having previously shown that our approach reproduces the constituent counting rule found in QCD, we now observe qualitative agreement between our predictions and the extracted data in the high-energy fixed-angle regime.  We also provide predictions for all other 22-to-22 pion scattering processes. Our approach can be extended to a broader range of meson and glueball scattering processes in various holographic QCD models.
YITP-25-192
@@ 1 Introduction
Since the introduction of holography Maldacena:1997re, growing evidence now suggests that string theory may provide a dual description of QCD in the low-energy regime (for pioneering results, see references Witten:1998zw; Karch:2002sh; Sakai:2004cn; for a selection of reviews, see references Aharony:1999ti; Kim:2012ey; Ammon:2015wua).  Originally, however, string theory emerged from efforts to study high-energy meson scattering.  In this context, the Veneziano formula Veneziano:1968yb achieved partial success in the Regge regime Green:1987sp.  Going beyond the Regge regime to the high-energy fixed-angle regime, dimensional counting arguments in QCD predict a leading-order power-law scaling in the Mandelstam variable $s$ for meson scattering amplitudes $\mathcal{A}$, as captured by the constituent counting rule Matveev:1973ra; Brodsky:1973kr; Brodsky:1974vy:
@@
where $m$ is the minimum number of hard constituents involved in the scattering, and $f(\theta_{i})$ encodes the dependence on the scattering angles $\theta_{i}$.  The constituent counting rule (1) appears to be consistent with all experimental studies to date when applied appropriately. For recent discussion, see reference Reed:2020dpf and references therein.  By contrast, string theory in Minkowski space demonstrates exponential rather than power-law suppression in $s$ in the high-energy fixed-angle regime.
@@
To revisit this situation in the context of holography, Polchinski and Strassler proposed studying scalar glueball scattering by analysing the dual dilaton superstring scattering in five-dimensional AdS space with an IR cut-off Polchinski:2001tt.  Using this phenomenological approach, which models QCD’s UV behaviour via a strongly coupled CFT, Polchinski and Strassler recovered the constituent counting rule.  More recently, in collaboration with Sugimoto Armoni:2024gqv, we applied their proposal to pion and $\rho$-meson scattering in a related holographic model, also recovering the constituent counting rule.  However, for processes involving $\rho$-mesons, we had to generalise the original proposal Polchinski:2001tt to reproduce the constituent counting rule.
@@
In this paper, we compare the predictions of our model Armoni:2024gqv, reformulated in the context of the hard-wall model, for the angular dependence of $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering with experimental data indirectly extracted from $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements reported in reference Bromberg:1983he.  We extract the data using two different models, the simpler of which assumes only a Yukawa-type nucleon-nucleon interaction mediated by the pion (details will be provided in Section 3).  The extracted data are thus off-shell. However, we expect deviations from on-shell behaviour to be qualitatively negligible when the Mandelstam variables in the underlying pion process far exceed the pion mass Martin:1976mb; Stampke:1982cf.
@@
Regarding our holographic model from reference Armoni:2024gqv, note that it is formulated within a five-dimensional asymptotically AdS (AAdS) background consistent with confinement and chiral symmetry breaking, hereafter referred to as AAdS space.  Our predictions in the limit of $s\to\infty$ are based on an ansatz for tree-level superstring scattering to leading order in $\alpha^{\prime}/R_{5}^{2}$ in this AAdS space (where $\alpha^{\prime}$ is the Regge slope and $R_{5}$ is the AdS radius), and rely only on its AdS region, which governs the leading $s$-dependence in the high-energy fixed-angle regime. However, in order to compare with experimental data at finite energies, the asymptotics are not sufficient and one must specify the background metric also in the IR region. For this purpose we choose the simplest confining AAdS background, which is AdS space with a finite IR cutoff, namely the hard-wall model also used in reference Polchinski:2001tt.
@@
Despite the limitations of our approach, we will show that, when combined with input from phenomenology, its predictions qualitatively agree with the data from reference Bromberg:1983he in the high-energy fixed-angle regime.  In contrast, agreement in the Regge regime is lacking, which could have been anticipated, as will be discussed in Section 2.2.
@@
This paper is organised as follows.  In Section 2, we review our holographic model and ansatz for pion scattering amplitudes from reference Armoni:2024gqv.  In Section 3, we present two models for indirectly extracting $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering data from $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements. In Section 4, we compare our predictions with data from reference Bromberg:1983he and offer predictions for all other 22-to-22 pion scattering processes.  Finally, in Section 5, we outline our findings and directions for future research.  Two appendices are also included.  Appendix A provides further comparisons to data for the discussion in Section 4.1, while Appendix B presents evidence suggesting that the $\pi^{-}p\to\pi^{+}\pi^{-}n$ data in reference Bromberg:1983he include a significant contribution from an underlying 4-point scattering process.
@@ 2 High-Energy Pion Scattering in Holography
This section begins with a review of our holographic model from reference Armoni:2024gqv in Section 2.1. In Section 2.2, we describe how the Polchinski–Strassler proposal Polchinski:2001tt can be used to write an ansatz for pion scattering amplitudes consistent with the constituent counting rule (1), following references Bianchi:2021sug; Armoni:2024gqv.  We also discuss the behaviour of this ansatz in the Regge regime. In Section 2.3, we precisely formulate our 4-point pion scattering amplitudes, addressing their regularisation and numerical evaluation.
@@ 2.1 Holographic QCD Model
The setup described above is consistent with confinement and chiral symmetry breaking. [^1] For a top-down example of a similar background in string theory, see reference Kuperstein:2008cq.
@@
As in reference Armoni:2024gqv, $B_{\mu}^{(n)}$ correspond to a tower of massive meson fields alternating between vector and axial-vector states, with the lightest state $B_{\mu}^{(1)}$ identified as the $\rho$-meson field. The zero mode $\varphi^{(0)}$ is identified as the massless pion field, while the remaining fields $\varphi^{(n)}$ for $n\geq 1$ correspond to the would-be Nambu–Goldstone bosons absorbed by $B_{\mu}^{(n)}$.
@@ 2.2 Scattering Amplitude Ansatz
The action (2.1) should be interpreted as the kinetic term for massless open strings, which forms part of the effective low-energy theory describing our mesons.  Meanwhile, computing high-energy pion scattering amplitudes in our model is expected to involve the exchange of the infinite tower of string states that UV-complete the action (2.1).  In a related setting, Polchinski and Strassler Polchinski:2001tt introduced an ansatz for dilaton scattering amplitudes that accounts for the complete string tower and reproduces the constituent counting rule (1) in a five-dimensional AdS space with an IR cut-off (for motivation behind their proposal, see also reference Brower:2006ea, particularly Section 2; for an explicit example of their proposal applied to a simple toy model in field theory, see Section 3.3 of reference Armoni:2024gqv).
Generalising the original proposal Polchinski:2001tt, reference Armoni:2024gqv proposed an ansatz in the context of the holographic model from Section 2.1 for pion and $\rho$-meson scattering amplitudes in the high-energy fixed-angle limit, with the ansatz for 4-point pion scattering given by:
@@
    In reference Armoni:2024gqv, following reference Polchinski:2001tt, it was shown that $p_{w}^{(i)}$ contribute as subleading corrections in $\alpha^{\prime}/R_{5}^{2}$ to the 4-point pion scattering amplitude $\mathcal{A}$ in the high-energy fixed-angle limit.  We will therefore neglect them in our analysis. [^3] Note that, for $\rho$-meson scattering, completely neglecting the components $p_{w}^{(i)}$ leads to a vanishing result. See reference Armoni:2024gqv for further details.
@@
Following reference Armoni:2024gqv, we now review how the pion scattering amplitude $\mathcal{A}$ given by equation (15) reproduces the constituent counting rule, focusing on tree-level scattering. To show this, it suffices to consider only the UV-region of the AAdS metric (5), and while the explicit form of the superstring scattering amplitude $\mathcal{S}$ in Minkowski space is also not required, it will be useful for our later analysis to consider the following expression Green:1987sp:
@@
where $g_{s}$ is the string coupling constant, $\mathrm{K}$ is the kinematical factor (see equation (7.4.42) in reference Green:1987sp), and the Chan–Paton factors, temporarily omitted above, will be restored in Section 2.3.
@@
Finally, the leading $s$-dependence of $\mathcal{A}$ in the high-energy fixed-angle limit is dictated only by the UV-region ($w\to\pm\infty$) of the AAdS metric (5), since all other contributions are suppressed by the exponential softness of the superstring scattering amplitude $\widetilde{\mathcal{S}}$ Polchinski:2001tt; Armoni:2024gqv. In this limit, the $s$-scaling of $\mathcal{A}$ agrees with the constituent counting rule (1):
@@
As discussed above and in reference Armoni:2024gqv, the leading $s$-dependence of the pion scattering amplitude $\mathcal{A}$ in the high-energy fixed-angle limit is determined only by the UV-region of the metric (5).  Consequently, the angular dependence associated with this leading-order behaviour should also be determined solely by the UV-region.  To extract this dependence, we adopt a simple approach, as already mentioned above.  We introduce an IR cut-off in the integral (15) and use only the AdS metric.  For this approach to be valid, the results should be independent of the IR cut-off, provided it is finite and non-zero.  As we will show in Section 2.3.1 based on numerical calculations, in the strict high-energy fixed-angle limit they are indeed independent of it.
@@
Now, to better inform our comparisons in Section 4.1, we will examine the behaviour of the pion scattering amplitude $\mathcal{A}$ outside the high-energy fixed-angle regime in the context of the hard-wall model, while briefly commenting on why these results do not constitute genuine predictions of our holographic model from references Armoni:2024gqv.
In QCD, the well-known Regge regime of 4-point hadronic scattering amplitudes is characterised by $|t|\lesssim\Lambda^{2}\ll s$, where $\Lambda$ is the QCD scale.  In this regime, string theory scattering amplitudes in Minkowski space show Regge behaviour $s^{\alpha(t)}$ (where $\alpha(t)$ denotes a linear function of $t$), which is broadly consistent with experiments Green:1987sp.  There also exists a lesser-known regime, $\Lambda^{2}\ll|t|\ll s$, in which string scattering amplitudes continue to exhibit Regge behaviour, whereas experimental data do not support such behaviour. In the following, we will restrict our attention to the broad regime $|t|/s\ll 1$, within which we will identify two distinct sub-regimes depending on the values of $t$ and $s$.  Our analysis closely draws on the related discussion from reference Polchinski:2001tt; Brower:2006ea.
@@
The function $f(\nu)\to 0$ as $\nu\to 0$ and $\nu\to\infty$, and attains an extremum at $\nu_{*}=2/\ln(s/|t|)$, while the function $g(\nu)\to\infty$ as $\nu\to 0^{+}$ and decreases monotonically to zero as $\nu\to\infty$. Finally, the entire integrand in equation (25) tends to zero both as $\nu\to 0$ and $\nu\to\infty$, and posses an extremum that is essentially set by the location of the extremum of $f(\nu)$. [^4] Strictly speaking, notice that the limit $\nu\to\infty$ lies outside the integration domain in equation (25). However, since we will later make use of the saddle point analysis, it is worthwhile to note that the integrand vanishes in this limit, allowing us to proceed with the standard analysis. A more precise evaluation shows that the extremum shifts slightly, although it remains of the same order, with the leading approximation giving $\nu_{*}\sim 1/\ln(s/|t|)$.  Moreover, it can be shown that this extremum is a local maximum.
@@
where we omitted all $s$-independent factors in the leading-order term and neglected subleading terms in $1/\ln(s/|t|)$.  In agreement with reference Polchinski:2001tt, we recover Regge behaviour in the regime where $\nu_{0}<\nu_{*}$, up to a logarithmic factor.  In our case, $\mathcal{A}\sim s^{\alpha(t)}$ with $\alpha(t)=1+\alpha^{\prime}tR_{5}^{2}/w_{0}^{2}$, where $\alpha^{\prime}R_{5}^{2}/w_{0}^{2}$ can be identified as the effective Regge slope.
@@
The results in equations (27) and (28) are consistent with the corresponding results of reference Polchinski:2001tt for scalar glueball scattering amplitudes, up to logarithmic factors. Now, in the context of the hard-wall model, the pion scattering amplitude $\mathcal{A}$ exhibits Regge behaviour at fixed, large $s$ and lower $|t|$ (as set by $\nu_{0}<\nu_{*}$), while this behaviour is lost at higher $|t|$. As noted in reference Polchinski:2001tt and references therein, the transition between these regimes is likely affected by multiloop effects. Given this, our tree-level approach is unlikely to reliably describe the Regge regime in a broad sense ($|t|/s\ll 1$) without distinguishing between the two regimes.
It is also worth pointing out that, following reference Brower:2002er, we do not expect that neglecting the components $p_{w}^{(i)}$ throughout the entire broad regime $|t|/s\ll 1$ is valid.
Now, let us also explain why these results likely do not follow through as genuine predictions of our holographic model from reference Armoni:2024gqv. The issue is that the analysis above relies solely on the AdS metric, and we have seen that our results depend on features such as the IR cut-off in the hard-wall AdS model. In this sense, it would be inconsistent to consider only the UV region of metric (5); if the leading behaviour depends on the IR cut-off, the behaviour of the metric away from the UV region must also be taken into account.

### v2

@@ Abstract
Following Polchinski and Strassler Polchinski and Strassler (2002) and our previous work Armoni et al. (2025), we study high-energy pion scattering in the holographic QCD hard-wall model.  In particular, we focus on comparing our predictions for the angular dependence of $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering with experimental data extracted from the process $\pi^{-}p\to\pi^{+}\pi^{-}n$.  Having previously shown that our approach reproduces the constituent counting rule found in QCD, we now observe qualitative agreement between our predictions and the extracted data in the high-energy fixed-angle regime.  We also provide predictions for all other 22-to-22 pion scattering processes. Our approach can be extended to a broader range of meson and glueball scattering processes in various holographic QCD models.
@@ 1 Introduction
Since the introduction of holography Maldacena (1998), growing evidence now suggests that string theory may provide a dual description of QCD in the low-energy regime (for pioneering results, see references Witten (1998); Karch and Katz (2002); Sakai and Sugimoto (2005); for a selection of reviews, see references Aharony et al. (2000); Kim et al. (2013); Ammon and Erdmenger (2015)).  Originally, however, string theory emerged from efforts to study high-energy meson scattering.  In this context, the Veneziano formula Veneziano (1968) achieved partial success in the Regge regime Green et al. (1988).  Going beyond the Regge regime to the high-energy fixed-angle regime, dimensional counting arguments in QCD predict a leading-order power-law scaling in the Mandelstam variable $s$ for meson scattering amplitudes $\mathcal{A}$, as captured by the constituent counting rule Matveev et al. (1973); Brodsky and Farrar (1973, 1975):
@@
where $m$ is the minimum number of hard constituents involved in the scattering, and $f(\theta_{i})$ encodes the dependence on the scattering angles $\theta_{i}$.  The constituent counting rule (1) appears to be consistent with all experimental studies to date when applied appropriately. For recent discussion, see reference Reed et al. (2021) and references therein.  By contrast, string theory in Minkowski space demonstrates exponential rather than power-law suppression in $s$ in the high-energy fixed-angle regime.
@@
To revisit this situation in the context of holography, Polchinski and Strassler proposed studying scalar glueball scattering by analysing the dual dilaton superstring scattering in five-dimensional AdS space with an IR cut-off Polchinski and Strassler (2002).  Using this phenomenological approach, which models QCD’s UV behaviour via a strongly coupled CFT, Polchinski and Strassler recovered the constituent counting rule.  More recently, in collaboration with Sugimoto Armoni et al. (2025), we applied their proposal to pion and $\rho$-meson scattering in a related holographic model, also recovering the constituent counting rule.  However, for processes involving $\rho$-mesons, we had to generalise the original proposal Polchinski and Strassler (2002) to reproduce the constituent counting rule.
@@
In this paper, we compare the predictions of our model Armoni et al. (2025), reformulated in the context of the hard-wall model, for the angular dependence of $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering with experimental data. Since there is no direct data for $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering, we use data which is indirectly extracted from $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements reported in reference Bromberg and others (1984a).  We extract the data using two different models, the simpler of which assumes only a Yukawa-type nucleon-nucleon interaction mediated by the pion (details will be provided in Section 3).  The extracted data are thus off-shell. However, we expect deviations from on-shell behaviour to be qualitatively negligible when the Mandelstam variables in the underlying pion process far exceed the pion mass Martin et al. (1976); Stampke (1982).
@@
Regarding our holographic model from reference Armoni et al. (2025), note that it is formulated within a five-dimensional asymptotically AdS (AAdS) background consistent with confinement and chiral symmetry breaking, hereafter referred to as AAdS space.  Our predictions in the limit of $s\to\infty$ are based on an ansatz for tree-level superstring scattering to leading order in $\alpha^{\prime}/R_{5}^{2}$ in this AAdS space (where $\alpha^{\prime}$ is the Regge slope and $R_{5}$ is the AdS radius), and rely only on its AdS region, which governs the leading $s$-dependence in the high-energy fixed-angle regime. However, in order to compare with experimental data at finite energies, the asymptotics are not sufficient and one must specify the background metric also in the IR region. For this purpose we choose the simplest confining AAdS background, which is AdS space with a finite IR cutoff, namely the hard-wall model also used in reference Polchinski and Strassler (2002). Note that as in the Sakai–Sugimoto model Sakai and Sugimoto (2005) mesons are incorporated without the need to modify the background, because in the quenched (large-NN) approximation mesons do not back react.
@@
Despite the limitations of our approach, we will show that, when combined with input from phenomenology, its predictions qualitatively agree with the data from reference Bromberg and others (1984a) in the high-energy fixed-angle regime.  In contrast, agreement in the Regge regime is lacking, which could have been anticipated, as will be discussed in Section $2.$
@@
This paper is organised as follows.  In Section 2, we review our holographic model and ansatz for pion scattering amplitudes from reference Armoni et al. (2025).  In Section 3, we present two models for indirectly extracting $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering data from $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements. In Section 4, we compare our predictions with data from reference Bromberg and others (1984a) and offer predictions for all other 22-to-22 pion scattering processes.  Finally, in Section 5, we outline our findings and directions for future research.  Two appendices are also included.  Appendix A provides further comparisons to data for the discussion in Section 4.1, while Appendix B presents evidence suggesting that the $\pi^{-}p\to\pi^{+}\pi^{-}n$ data in reference Bromberg and others (1984a) include a significant contribution from an underlying 4-point scattering process.
@@ 2 High-Energy Pion Scattering in Holography
This section begins with a review of our holographic model from reference Armoni et al. (2025) in Section 2.1. In Section 2.2, we describe how the Polchinski–Strassler proposal Polchinski and Strassler (2002) can be used to write an ansatz for pion scattering amplitudes consistent with the constituent counting rule (1), following references Bianchi et al. (2022); Armoni et al. (2025).  We also discuss the behaviour of this ansatz in the Regge regime. In Section 2.3, we precisely formulate our 4-point pion scattering amplitudes, addressing their regularisation and numerical evaluation.
@@ 2.1 Holographic QCD Model
The setup described above is consistent with confinement and chiral symmetry breaking. [^1] For a top-down example of a similar background in string theory, see reference Kuperstein and Sonnenschein (2008).
@@
As in reference Armoni et al. (2025), $B_{\mu}^{(n)}$ correspond to a tower of massive meson fields alternating between vector and axial-vector states, with the lightest state $B_{\mu}^{(1)}$ identified as the $\rho$-meson field. The zero mode $\varphi^{(0)}$ is identified as the massless pion field. It has the correct quantum numbers and is massless because of its role as the Nambu–Goldstone boson of chiral symmetry breaking. The remaining fields $\varphi^{(n)}$ for $n\geq 1$ correspond to the would-be Nambu–Goldstone bosons absorbed by $B_{\mu}^{(n)}$.
@@ 2.2 Scattering Amplitude Ansatz
The action (2.1) should be interpreted as the kinetic term for massless open strings, which forms part of the effective low-energy theory describing our mesons.  Meanwhile, computing high-energy pion scattering amplitudes in our model is expected to involve the exchange of the infinite tower of string states that UV-complete the action (2.1).  In a related setting, Polchinski and Strassler Polchinski and Strassler (2002) introduced an ansatz for dilaton scattering amplitudes that accounts for the complete string tower and reproduces the constituent counting rule (1) in a five-dimensional AdS space with an IR cut-off (for motivation behind their proposal, see also reference Brower et al. (2007), particularly Section 2; for an explicit example of their proposal applied to a simple toy model in field theory, see Section 3.3 of reference Armoni et al. (2025)).
Generalising the original proposal Polchinski and Strassler (2002), reference Armoni et al. (2025) proposed an ansatz in the context of the holographic model from Section 2.1 for pion and $\rho$-meson scattering amplitudes in the high-energy fixed-angle limit, with the ansatz for 4-point pion scattering given by:
@@
    In reference Armoni et al. (2025), following reference Polchinski and Strassler (2002), it was shown that $p_{w}^{(i)}$ contribute as subleading corrections in $\alpha^{\prime}/R_{5}^{2}$ to the 4-point pion scattering amplitude $\mathcal{A}$ in the high-energy fixed-angle limit.  We will therefore neglect them in our analysis. [^3] Note that, for $\rho$-meson scattering, completely neglecting the components $p_{w}^{(i)}$ leads to a vanishing result. See reference Armoni et al. (2025) for further details.
@@
Following reference Armoni et al. (2025), we now review how the pion scattering amplitude $\mathcal{A}$ given by equation (15) reproduces the constituent counting rule, focusing on tree-level scattering. To show this, it suffices to consider only the UV-region of the AAdS metric (5), and while the explicit form of the superstring scattering amplitude $\mathcal{S}$ in Minkowski space is also not required, it will be useful for our later analysis to consider the following expression Green et al. (1988):
@@
where $g_{s}$ is the string coupling constant, $\mathrm{K}$ is the kinematical factor (see equation (7.4.42) in reference Green et al. (1988)), and the Chan–Paton factors, temporarily omitted above, will be restored in Section 2.3.
@@
Finally, the leading $s$-dependence of $\mathcal{A}$ in the high-energy fixed-angle limit is dictated only by the UV-region ($w\to\pm\infty$) of the AAdS metric (5), since all other contributions are suppressed by the exponential softness of the superstring scattering amplitude $\widetilde{\mathcal{S}}$ Polchinski and Strassler (2002); Armoni et al. (2025). In this limit, the $s$-scaling of $\mathcal{A}$ agrees with the constituent counting rule (1):
@@
As discussed above and in reference Armoni et al. (2025), the leading $s$-dependence of the pion scattering amplitude $\mathcal{A}$ in the high-energy fixed-angle limit is determined only by the UV-region of the metric (5).  Consequently, the angular dependence associated with this leading-order behaviour should also be determined solely by the UV-region.  To extract this dependence, we adopt a simple approach, as already mentioned above.  We introduce an IR cut-off in the integral (15) and use only the AdS metric.  For this approach to be valid, the results should be independent of the IR cut-off, provided it is finite and non-zero.  As we will show in Section 2.3.1 based on numerical calculations, in the strict high-energy fixed-angle limit they are indeed independent of it.
@@
Now, to better inform our comparisons in Section 4.1, we will examine the behaviour of the pion scattering amplitude $\mathcal{A}$ outside the high-energy fixed-angle regime in the context of the hard-wall model, while briefly commenting on why these results do not constitute genuine predictions of our holographic model from reference Armoni et al. (2025).
In QCD, the well-known Regge regime of 4-point hadronic scattering amplitudes is characterised by $|t|\lesssim\Lambda^{2}\ll s$, where $\Lambda$ is the QCD scale.  In this regime, string theory scattering amplitudes in Minkowski space show Regge behaviour $s^{\alpha(t)}$ (where $\alpha(t)$ denotes a linear function of $t$), which is broadly consistent with experiments Green et al. (1988).  There also exists a lesser-known regime, $\Lambda^{2}\ll|t|\ll s$, in which string scattering amplitudes continue to exhibit Regge behaviour, whereas experimental data do not support such behaviour. In the following, we will restrict our attention to the broad regime $|t|/s\ll 1$, within which we will identify two distinct sub-regimes depending on the values of $t$ and $s$.  Our analysis closely draws on the related discussion from reference Polchinski and Strassler (2002); Brower et al. (2007).
@@
The function $f(\nu)\to 0$ as $\nu\to 0$ and $\nu\to\infty$, and attains an extremum at $\nu_{*}=2/\ln(s/|t|)$, while the function $g(\nu)\to\infty$ as $\nu\to 0^{+}$ and decreases monotonically to zero as $\nu\to\infty$. Finally, the entire integrand in equation (25) tends to zero both as $\nu\to 0$ and $\nu\to\infty$, and possesses an extremum that is essentially set by the location of the extremum of $f(\nu)$. [^4] Strictly speaking, notice that the limit $\nu\to\infty$ lies outside the integration domain in equation (25). However, since we will later make use of the saddle point analysis, it is worthwhile to note that the integrand vanishes in this limit, allowing us to proceed with the standard analysis. A more precise evaluation shows that the extremum shifts slightly, although it remains of the same order, with the leading approximation giving $\nu_{*}\sim 1/\ln(s/|t|)$.  Moreover, it can be shown that this extremum is a local maximum.
@@
where we omitted all $s$-independent factors in the leading-order term and neglected subleading terms in $1/\ln(s/|t|)$.  In agreement with reference Polchinski and Strassler (2002), we recover Regge behaviour in the regime where $\nu_{0}<\nu_{*}$, up to a logarithmic factor.  In our case, $\mathcal{A}\sim s^{\alpha(t)}$ with $\alpha(t)=1+\alpha^{\prime}tR_{5}^{2}/w_{0}^{2}$, where $\alpha^{\prime}R_{5}^{2}/w_{0}^{2}$ can be identified as the effective Regge slope. Note that the intercept we obtain differs from the phenomenologically accepted value, hence we do not anticipate a good agreement with the data at small angles.
@@
The results in equations (27) and (28) are consistent with the corresponding results of reference Polchinski and Strassler (2002) for scalar glueball scattering amplitudes, up to logarithmic factors. Now, in the context of the hard-wall model, the pion scattering amplitude $\mathcal{A}$ exhibits Regge behaviour at fixed, large $s$ and lower $|t|$ (as set by $\nu_{0}<\nu_{*}$), while this behaviour is lost at higher $|t|$. As noted in reference Polchinski and Strassler (2002) and references therein, the transition between these regimes is likely affected by multiloop effects. Given this, our tree-level approach is unlikely to reliably describe the Regge regime in a broad sense ($|t|/s\ll 1$) without distinguishing between the two regimes.
It is also worth pointing out that, following reference Brower et al. (2007), we do not expect that neglecting the components $p_{w}^{(i)}$ throughout the entire broad regime $|t|/s\ll 1$ is valid.
Now, let us also explain why these results likely do not follow through as genuine predictions of our holographic model from reference Armoni et al. (2025). The issue is that the analysis above relies solely on the AdS metric, and we have seen that our results depend on features such as the IR cut-off in the hard-wall AdS model. In this sense, it would be inconsistent to consider only the UV region of metric (5); if the leading behaviour depends on the IR cut-off, the behaviour of the metric away from the UV region must also be taken into account.